OpenStudy (anonymous):

help? 2 questions c: medal!

OpenStudy (anonymous):

3. What is the value of x to the nearest tenth? a.18 b.19 c.71 d.72

OpenStudy (mertsj):

In relation to the angle x, what sides do you know: hypotenuse, adjacent, opposite. Choose.

OpenStudy (anonymous):

hypotenuse?

OpenStudy (mertsj):

Do you understand that the hypotenuse is always the longest sides in a right triangle and that it is opposite the right angle?

OpenStudy (mertsj):



OpenStudy (anonymous):

yes I understand

OpenStudy (mertsj):

So I repeat my original question: In relation to the angle x, what two sides do you know?

OpenStudy (mertsj):

dang, buddy, the triangle only has 3 sides. If it's not the hypotenuse, what is left?

OpenStudy (anonymous):

I know sorry my screen froze I guess adjacent and opposite

OpenStudy (mertsj):

Which is the adjacent side: 11 or 33? Remember adjacent means "next to"

OpenStudy (anonymous):

33

OpenStudy (mertsj):

Now, you know the adjacent side and the opposite side, what trig function does that make you think of?

OpenStudy (anonymous):

tan right?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

\[\tan x=\frac{opp}{adj}\]

OpenStudy (mertsj):

Fill in the sides and solve the equation for x

OpenStudy (anonymous):

I got 0.00581 ? ...

OpenStudy (mertsj):

What did you get for the tangent?

OpenStudy (anonymous):

that's what I got I put tan 11/33

OpenStudy (mertsj):

\[\tan x=\frac{11}{33}=\frac{1}{3}=.3333333...\]

OpenStudy (anonymous):

how did you get the last part?

OpenStudy (mertsj):

What do you think 1 divided by 3 is?

OpenStudy (anonymous):

and the answers are a.18 b.19 c.71 d.72 /:

OpenStudy (anonymous):

ohhh ok nvm >_<

OpenStudy (mertsj):

So type .33333333 into your calculator and then push the inverse tangent button

OpenStudy (mertsj):

Or, depending on your calculator push inverse tangent button and then .33333333 =

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

okay im using this calculator because I don't have one and still doing that I get the same answer .. http://web2.0calc.com/

OpenStudy (mertsj):

Push 2nd Then the tan button will say atan which stands for "arctangent" Then enter .3333333333 Then push =

OpenStudy (anonymous):

18.4 ! a

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw